OpenStudy (anonymous):

If a, b ,c are real numbers such that |a-b|>=|c|, |b-c|>=|a|, |c-a|>=|b|, then prove that one of a, b, c is the sum of the other two.

OpenStudy (shubhamsrg):

*

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

what does * mean?

ganeshie8 (ganeshie8):

just it gives us a reminder when somebody updates this post... :)

OpenStudy (experimentx):

put a = b + c |b-c|>=|a| = |b + c|, this is only true if c<=0, b>=0

OpenStudy (shubhamsrg):

we are given a,b,c are real numbers, hence we can plot them on the x axes. Now due to symmetrical behavior of a,b and c, let can assume a<b<c here, I am gonna use the fact that |x-y| means distance between x and y. now, |a-b| >=|c| |b-c|>=|a| adding both, |a-b| + |b-c| >= |c|+|a| now since a<b<c, |a-b| + |b-c| = |c-a| hence |c-a| >= |c| + |a|, note that greater is never possible and equal is only possible when a and c have opposite signs ! since a<c, then a must be negative and c must be positive,and b lies somewhere between a and c now, |a-b| >= |c| |a-b| + |a| >= |c| + |a| |a-b| + |a| >= |c-a| |a| >= |c-a| - |a-b| |a| >= |c-b| but we are given |a| <= |c-b| this means |a| = |c-b| its easy to force the final conclusion from here.

OpenStudy (anonymous):

Thanks @shubhamsrg. I am going to justify why |c-a| - |a-b| = |c-b| in this problem and then finish it from where you left off. |c-a| = the distance between a and c (call this AC) |a-b| = the distance between a and b (call this AB) |c-b| = the distance between b and c (call this BC) Since we assume a<b<c, AC - AB = BC, which means |c-a| - |a-b| = |c-b| is true. Skip ahead to |a| = |c-b|, which can be rewritten as |a| = c-b, since b<c means |c-b| = c-b. |a| = |c - b| |a| = c - b |a| + b = c Since a is negative, |a| = -a. |a| + b = c -a + b = c b = c + a Or, a+c=b.

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

 if we multiply all the inequalities 1 , 2 and 3,